# Exploring Lattice Thermal Conductivity Models via Interpretable Deep Learning to Accelerate Materials Discovery

### v8

@@ III DISCUSSION
In the past, researchers often relied on empirical or semi-empirical models to calculate the $\kappa_{\rm L}$ of materials. However, as the range of explored materials continues to expand, the inaccuracies inherent in empirical models have become increasingly problematic. Machine learning-based $\kappa_{\rm L}$ modeling has achieved remarkable accuracy, but inprovments in accuracy alone offers limited contributions to the advancement of materials science. Black-box models, such as neural networks and ensemble learning, excel in accuracy but their complex structures hinder the understanding of feature-target relationships. In contrast, white-box models like symbolic regression provide transparent structures, making their internal mechanisms interpretable, but this often comes at the cost of accuracy. Traditionally, the interpretability and accuracy of ML or DL models have been seen as mutually exclusive, akin to “having your cake and eating it too.” However, in the context of $\kappa_{\rm L}$ modeling, interpretable DL models like KAN have demonstrated significant success in balancing both accuracy and interpretability.
In this study, we employed KAN to model $\kappa_{\rm L}$, demonstrating that its performance in terms of both accuracy and robustness is fully comparable to that of black-box models. For the interpretability analysis of white-box models, we considered feature interaction effects. Sensitivity analysis results based on the KAN and SISSO models revealed that KAN can accurately extract key features, highlighting its superior interpretability in this context. Additionally, the analytical expressions derived from KAN serve as a valuable supplement to traditional empirical models. We also developed a two-stage high-throughput prediction framework, integrating CGCNN to accurately predict the primary physical features required for $\kappa_{\rm L}$. For $C_{\text{v}}$, a property that CGCNN cannot predict accurately, DFT calculations were used. Building on this foundation, the predictions from the KAN pretraining model aligned closely with MD validation results, enabling us to successfully identify three new materials, one of which was confirmed as an excellent thermal conductor. Compared to using MD to directly compute $\kappa_{\rm L}$ for unknown materials, our high-throughput prediction framework requires only DFT calculations of $C_{\text{v}}$ to achieve MD-level predictive accuracy. This work not only accelerates the prediction of material thermal conductivity but also enhances our understanding of the underlying physical mechanisms. By balancing accuracy and interpretability, KAN demonstrates its potential as a powerful tool for advancing materials informatics.
@@ IV.1 Kolmogorov-Arnold Networks (KANs)
$\phi\left(x\right)=\omega\left[b\left(x\right)+\text{spline}\left(x\right)\right]=\omega\left[\frac{x}{1+{{\text{e}}^{-x}}}+\sum\limits_{i}{{{c}_{i}}{{B}_{i}}\left(x\right)}\right]$ (10)
@@ IV.6 Implementation of DFT & GPUMD methods
For the calculation of $\kappa_{\rm L}$, both the phonon Boltzmann Transport Equation (BTE) method and Molecular Dynamics (MD) simulations are used. For the stable Be2C structure, specific heat and $\kappa_{\rm L}$ are computed using ShengBTE [100]. The force constants are generated with Hiphive [101]. For other structures, specific heat is calculated using Phonopy [102], while $\kappa_{\rm L}$ is determined through GPUMD [103]. AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ns. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,\text{\AA}^{-1}$ has been tested and found to be optimal for this study.

### v9

@@ III DISCUSSION
In the past, researchers often relied on empirical or semi-empirical models to calculate the $\kappa_{\rm L}$ of materials. However, as the range of explored materials continues to expand, the inaccuracies inherent in empirical models have become increasingly problematic. Machine learning-based $\kappa_{\rm L}$ modeling has achieved remarkable accuracy, but improvements in accuracy alone offers limited contributions to the advancement of materials science. Black-box models, such as neural networks and ensemble learning, excel in accuracy but their complex structures hinder the understanding of feature-target relationships. In contrast, white-box models like symbolic regression provide transparent structures, making their internal mechanisms interpretable, but this often comes at the cost of accuracy. Traditionally, the interpretability and accuracy of ML or DL models have been seen as mutually exclusive, akin to “having your cake and eating it too.” However, in the context of $\kappa_{\rm L}$ modeling, interpretable DL models like KAN have demonstrated significant success in balancing both accuracy and interpretability.
Our work employed KAN to model $\log(\kappa_{\rm L})$, demonstrating that its performance in terms of both accuracy and robustness is fully comparable to that of black-box models. For the interpretability analysis of white-box models, we considered feature interaction effects and dependencies. Sensitivity analysis results based on the KAN and SISSO models revealed that KAN can accurately extract key features of $\kappa_{\rm L}$, highlighting its superior interpretability in this context. We combined KAN with CGCNN to construct a two-stage prediction framework, where CGCNN predicts key features, and the KAN pretraining model maps these features to $\log(\kappa_{\rm L})$; for physical properties like $C_{\text{v}}$, which CGCNN struggles to predict accurately, we used DFT calculations to complement the framework. In the case of unlabeled new samples, we applied $k$-means clustering to identify 10 potential thermal conductors/insulators. Among them, the $\kappa_{\rm L}$ values of CN2, CrB4, Be2C, and KRbAu2 were verified by DFT and MD, yielding results highly consistent with deep learning predictions, which sufficiently demonstrate KAN’s excellent extrapolation capability. We visualized the phonon spectra, PDOS, -pCOHP, and other information of Be2C and KRbAu2 using DFT, analyzing the phonon transport mechanisms of both materials from the perspectives of chemical bonding, atomic mass, and structural stability, providing theoretical insights as reference.
@@ IV.1 Kolmogorov-Arnold Networks (KANs)
$\phi\left(x\right)$ $=\omega\left[b\left(x\right)+\text{spline}\left(x\right)\right]$
$=\omega\left[\frac{x}{1+{{\text{e}}^{-x}}}+\sum\limits_{i}{{{c}_{i}}{{B}_{i}}\left(x\right)}\right]$ (10)
@@ IV.6 Implementation of DFT & GPUMD methods
For the calculation of $\kappa_{\rm L}$, both the phonon BTE method and MD simulations are used. For the stable structures, $\kappa_{\rm L}$ are computed using phono3py [108, 109, 110]. While $\kappa_{\rm L}$ of unstable structures are determined through GPUMD [111] AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ns. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,{\rm\AA}^{-1}$ has been tested and found to be optimal for this study. Specific heat and phonon dispersions are calculated using Phonopy [112]. The pCOHP analysis is performed using the LOBSTER code [113]. Structural visualizations and ELF analysis are conducted using the VESTA software [114].